OpenStudy (baby456):

Prove algebraically that 2/x-3 +2 is equal to 2x-4/x-3

OpenStudy (mathmale):

For clarity, it is important that you enclose any denominators in parentheses: 2 / (x-3), NOT 2 / x-3, and -4 / (x-3), not -4 / x-3. To answer this question, combine the 2 terms on the left side and combine the 2 terms on the right side. Then compare your results.

OpenStudy (dayakar):

\[\frac{ 2 }{ x-3 }+2 =\frac{ 2 }{ x-3}+ \frac{ 2(x-3) }{ (x-3) }\] [ since x-3/x-3 =1] can u simplify it

OpenStudy (mathmale):

Yes, good suggestion. Simplify the right side of the equation\[\frac{ 2 }{ x-3 }+2 =\frac{ 2 }{ x-3}+ \frac{ 2(x-3) }{ (x-3) }\]

OpenStudy (baby456):

OK but my teacher means to prove they are the same thing

OpenStudy (baby456):

Like they should equal each other

OpenStudy (baby456):

@mathmale and @dayakar

OpenStudy (dayakar):

if u simplify above expression u can get the equal one

OpenStudy (mathmale):

I understand what your teacher is asking you to do. My point is that there is ambiguity in the way you have entered this equation, and I'm asking you to address (fix) that ambiguity.